OpenStudy (anonymous):

divide and simplify please help t^2+8t/t^2+5t-24 divided by t/t+3 = type a fraction

OpenStudy (anonymous):

factor it first \[t(t+8)/(t+8)(t-3) \div\ t/t+3\]= use the reciprocal \[t(t+8)/(t+8)(t-3) \times\ t=3/t\] then use the cancellation method. the answer will be.\[ t+3/t-3\]